Meiji Isetan Corp. of Japan has two regional divisions with headquarters in Osaka and Yokohama. Selected data on the two divisions follow: Division Osaka Yokohama Sales Net operating income 5 9,300,000 23,000,000 $ 651,000 2,070,000 Average operating assets s2,325,000 11,500,000 Required: . For each division, compute the return on investment (ROD in terms of margin and turnover 2. Assume that the company evaluates performance using residual income and that the minimum required rate of return for any division is 14%. Compute the residual income for each division. 3. Is Yokohamas greater amount of residual income an indication that it is better managed?
media%2F8c5%2F8c5dd9a9-851a-4000-b03c-84
media%2F2d9%2F2d949605-c493-4857-a341-cd
Answer #1

1.

Division
Osaka Yokohama
Sales (a) 9300000 23000000
Net operating income (b) 651000 2070000
Average operating assets (c) 2325000 11500000
Margin (d) = (b)/(a) 0.07 0.09
Turnover (e) = (a)/(c) 4 2
ROI (f) = (d) x (e) 28% 18%

2.

Division
Osaka Yokohama
Average operating assets (a) 2325000 11500000
Minimum required rate of return (b) 14% 14%
Minimum acceptable net operating income (c) = (a) x (b) 325500 1610000
Net operating income (d) 651000 2070000
Residual income (e) = (d) - (c) 325500 460000